# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v1

@@ 6.1.1. Personalization
Personalized responses present a double-edged sword—while they can enhance the adoption and utility of CataractBot by providing more relevant information (Yang et al., 2023), they may also raise privacy concerns among certain users, potentially reducing usage. Additionally, personalized responses may increase the workload for experts, as they would need to refer to patient-specific documents for verification purposes. Many of these verification edits by experts may not be suitable for inclusion in the KB due to there patient-specific nature. Given our demographic of individuals with low literacy and low tech-literacy, individuals might require assistance in making an informed decision about consenting to the use of their health data (Okolo et al., 2021). Also, adopting a binary approach–either sharing all medical records or none–may not align with user preferences. Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access, in line with the philosophy of the National Digital Health Mission (NDHM) (National Health Authority and Family Welfare, 2020) in India.
To achieve this, a potential solution could be to ask users to manually upload images/PDFs of relevant documents (such as consultation sheets, discharge summaries, and test reports) they wish the bot to access. This approach empowers users to control the information available to the chatbot for a more personalized response. Currently, CataractBot operates as a reactive, user-driven system. However, with personalization features, we propose expanding its role to drive interactions at crucial points during the patient’s treatment journey, providing advice on comorbidities and medication reminders.
6.1.2. Reduce Workload for Experts
@@
Deploying CataractBot in the hospital increased the workload for patient coordinators (due to onboarding and verification) and doctors (due to verification and KB updates). To alleviate their burden, we propose ways to streamline CataractBot-related tasks and to use CataractBot to support them in their day-to-day activities. First, to simplify patient onboarding, we propose installing QR codes with a concise CataractBot introduction at key locations in the hospital. Users can scan the code, self-onboard by providing necessary information (language preference, operating doctor, and surgery date), and instantly engage with the bot. Users should receive a brief tutorial video on interaction guidelines, emphasizing that the bot is not intended for emergencies. Second, CataractBot should be utilized to deliver general information about the surgery and provide status updates to patients and attendants, thereby assisting experts in managing their other responsibilities. Supporting routine procedures, as recommended by Li et al. (2023), would enable experts focus on more specific and meaningful conversations with their patients during their limited face-to-face time. Third, to reduce the verification load for experts, we propose a system to pre-verify frequently asked questions automatically based on experts’ repeated confirmations of correctness. Fourth, we recommend introducing a feature that allows experts to ‘pause’ CataractBot by specifying their ‘unavailable until ¡date time¿’ status in natural language. This not only helps prevent disturbances for experts during surgery days, outside working hours, or when on leave, but also ensures that information seekers are better informed of the experts’ availability, also suggested by Yang et al. (2023). Finally, regarding audio messages, the lack of non-verbal cues hindered experts’ ability to deliver care (similar to (Li et al., 2023)). To address the limitations of translation and transcription technologies, the information-seeker’s original audio clip and transcribed query should be shared with the experts.
@@ 6.1.3. Usefulness
Prior work found that patients often seek information to prepare for a doctor consultation, verify a diagnosis or treatment plan, and assess the need for clinical intervention (Attfield et al., 2006). However, access to CataractBot is currently provided post-scheduling of the surgery. Introducing the bot before surgery scheduling might conflict with hospital policies, as patients could potentially cancel surgeries after using the bot, resulting in resource wastage for both the bot and experts. For consultation-related information needs, a similar bot connected with the user’s family doctor could be enabled to address generic medical information needs. Based on our findings, we recommend keeping CataractBot enabled for a month post-surgery. During bot interactions, a few participants expressed frustration with IDK responses (similar to previous findings (Xiao et al., 2023)), leading to discontinuation of usage. To address this, we suggest enhancing IDK responses by providing general information about the topic, offering tentative recommendations with disclaimers, and guiding users to alternative information sources. Lastly, considering the visual capabilities of recent LLM models like GPT-4V (a multimodal model), CataractBot should be enhanced to handle image-based inputs, aligning with recommendations from previous work (Jain et al., 2018). For instance, patients could upload images of medications instead of providing written details. While this could improve the user experience, it may also increase the workload for experts.
6.1.4. Scaling
In this work, the use of CataractBot was limited to a small sample of 49 patients and attendants over 2 months, while Sankara Eye Hospital performs over 50 cataract surgeries daily. Scaling CataractBot to cater to all cataract patients would require several modifications. These include implementing a QR code-based self-onboarding system, establishing a more comprehensive custom knowledge base from the outset, streamlining the verification process involving multiple doctors and coordinators for each patient (considering the potential non-participation of some experts), and improving the KB update process by engaging multiple KB Update Experts. Even within our short deployment, we noticed the knowledge base becoming outdated and the KB Update expert adding incorrect information. For longitudinal deployments, we recommend engaging multiple hospital staff in managing the KB to reduce reliance on a single expert, and implementing a simple method for the staff to add new documents to the KB. Recognizing the impracticality of manual updates in rapidly changing environments (Xiao et al., 2023), we propose automatic updates for specific information sources, such as standard operating procedures.
Cataract surgery, being short, straightforward, commonplace, highly successful, and extensively documented, resulted in minimal queries from participants to CataractBot. A chatbot similar to CataractBot could provide more value to information seekers undergoing complex and less common treatments, like glaucoma surgery or radiation therapy for head and neck cancer. With the existing framework for constructing experts-in-the-loop bot, creating such bots should be feasible333We will open-source our codebase to help the research community build and deploy their own end-to-end expert bots.. Furthermore, the applicability of such a system need not be confined to the medical domain; other fields, such as legal or agriculture (Jain et al., 2018), could benefit from similar chatbots. In these scenarios, end-users (clients or farmers) could receive synchronous responses from an LLM operating on a custom knowledge base and subsequently obtain verified responses asynchronously from experts (lawyers or agricultural experts).
@@ 6.2. Limitations
We acknowledge several limitations of this work. First, the results should be interpreted considering the specificities of the study, including the relatively small sample size, users’ first encounter with a medical chatbot, and the cataract treatment scenario. The positive user responses and other conclusions should be validated by future, preferably longitudinal studies. Second, we need to conduct system-level analysis, such as scalability testing, power measurement, etc., in future work before scaling the proposed solution to the masses. Third, CataractBot uses responses from the expert-in-the-loop after approval by the KB Update expert to incrementally update the custom KB; however, the new information may conflict with the bot’s older KB. This issue needs to be resolved and requires further investigation. Finally, although participants showed high trust in the initial encounter, IDK responses and inaccurate information in daily usage can lead to distrust of the system. We also note the potential risk of disseminating inaccurate medical information at scale with chatbot technologies. Therefore, the knowledge base should be carefully developed, reviewed, and updated by experts.
@@ 7. Conclusion
Building a system that addresses patients’ information needs by providing expert-verified information is an open problem in healthcare. To tackle this, we propose a novel solution—a LLM-powered expert-in-the-loop chatbot framework, that utilizes RAG (Retrieval Augmented Generation) over a custom knowledge base to provide synchronous responses, and utilizes experts to provide verified responses asynchronously. Our in-the-wild study involving 49 participants (patients and attendants) and 6 experts (4 doctors and 2 coordinators) demonstrated the positive impact of this technological intervention. Patients not only trusted the chatbot’s response but were also willing to wait, understanding that their busy doctors were verifying them. Simultaneously, doctors and coordinators, despite their hectic schedules, squeezed time during breaks and at home to help patients by verifying LLM-generated responses. The favorable feedback from various stakeholders in the healthcare ecosystem indicates that a chatbot, delivered through ubiquitous smartphones and WhatsApp, can effectively enhance information access in critical healthcare settings, even for individuals with limited literacy and technology experience.
@@ Acknowledgements.
- Larasati et al. (2023) Retno Larasati, Anna De Liddo, and Enrico Motta. 2023. Meaningful Explanation Effect on User’s Trust in an AI Medical System: Designing Explanations for Non-Expert Users. ACM Trans. Interact. Intell. Syst. 13, 4, Article 30 (dec 2023), 39 pages. https://doi.org/10.1145/3631614
@@
- ter Hofstede et al. (1996) A. H. M. ter Hofstede, H. A. Proper, and Th. P. van der Weide. 1996. Query Formulation as an Information Retrieval Problem. Comput. J. 39, 4 (01 1996), 255–274. https://doi.org/10.1093/comjnl/39.4.255 arXiv:https://academic.oup.com/comjnl/article-pdf/39/4/255/1916704/390255.pdf
@@
- Yang et al. (2023) Ziqi Yang, Xuhai Xu, Bingsheng Yao, Shao Zhang, Ethan Rogers, Stephen Intille, Nawar Shara, Dakuo Wang, et al. 2023. Talk2Care: Facilitating Asynchronous Patient-Provider Communication with Large-Language-Model. arXiv preprint arXiv:2309.09357 (2023). arXiv:2309.09357 [cs.CL]

### v2

@@ 7.2.1. Personalization
Personalized responses present a double-edged sword—while they can enhance the adoption and utility of CataractBot by providing more relevant information, they may also raise privacy concerns among certain users, potentially reducing usage. Additionally, personalized responses could increase the workload of experts, as they would need to refer to specific patient documents for verification. Many of these verification edits by experts may not be suitable for inclusion in the KB due to their patient-specific nature. Given our demographic of low literate and low tech-literate users, they might require assistance in making informed decisions about consenting to the use of their health data (Okolo et al., 2021). Also, adopting a binary approach–either sharing all medical records or none–may not align with user preferences. Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access, in line with the philosophy of the National Digital Health Mission (NDHM) (National Health Authority and Family Welfare, 2020) in India. A solution could be to ask users to manually upload relevant documents (such as consultation sheets, discharge summaries, and test reports) or textually enter details they wish the bot to access. This approach can empower users to control the information available to the bot for a more personalized experience. Currently, CataractBot operates as a user-driven system. With personalization features, we propose expanding its role to drive interactions at crucial points during the patient’s journey, e.g., providing medication reminders.
7.2.2. Reduce Workload of Experts
@@
Deploying CataractBot in the hospital increased the workload of patient coordinators (due to onboarding and verification) and doctors (due to verification and KB updates). To alleviate their burden, we propose ways to streamline CataractBot-related tasks and to use CataractBot to support them in their day-to-day activities. First, to simplify patient onboarding, QR codes can be installed at key locations in the hospital. Users can scan the code, self-onboard by providing necessary information (such as language preference, operating doctor, and surgery date), and instantly engage with the bot. Users should receive a brief tutorial video on interaction guidelines, emphasizing that the bot is not intended for emergencies. Second, CataractBot should be utilized to deliver general information about the surgery and provide status updates to patients and attendants, thereby assisting experts in managing their other responsibilities. This would enable experts focus on more specific and meaningful conversations with their patients during their limited face-to-face time. Third, to reduce the verification load for experts, we propose a system to pre-verify frequently asked questions automatically based on experts’ repeated confirmations of correctness. Finally, regarding audio messages, the lack of non-verbal cues hindered experts’ ability to deliver care (similar to (Li et al., 2023)). To address the limitations of translation and transcription technologies, the information seeker’s original audio clip should also be shared with the experts.
@@ 7.2.3. Usefulness
Attfield et al. (2006) reported that patients seek information to prepare for a doctor consultation, verify a diagnosis or treatment plan, and assess the need for clinical intervention. However, access to CataractBot is currently provided only post-scheduling of the surgery. Introducing the bot before this might conflict with hospital policies, as patients could potentially cancel surgeries after using the bot, resulting in resource wastage for both the bot and experts. During bot interactions, a few participants expressed frustration with IDK responses (similar to previous findings (Xiao et al., 2023)), leading to discontinuation of usage. To address this, we suggest enhancing IDK responses by providing general information about the topic, offering tentative recommendations with disclaimers, and guiding users to alternative information sources. Lastly, considering the visual capabilities of recent LLM models like GPT-4V (a multimodal model), CataractBot should be enhanced to handle image-based inputs, aligning with recommendations from previous work (Jain et al., 2018). For instance, patients could upload images of medications instead of providing written details. While this could improve the user experience for information seekers, it may also increase the workload for experts.
@@ 7.3. Limitations
We acknowledge several limitations of this work. First, the results should be interpreted considering the specificities of the study, including the relatively small sample size, users’ first encounter with a medical chatbot, and the cataract treatment scenario. The positive user responses and other conclusions should be validated by future, preferably longitudinal studies. Second, system-level analyses, such as scalability testing and power measurement, are needed before scaling the proposed solution to the masses. Third, CataractBot uses expert responses to incrementally update the custom KB; however, the new information may conflict with the bot’s older KB. This issue requires further investigation. Finally, we note the potential risk of disseminating inaccurate medical information at scale with chatbot and LLM technologies. Therefore, stronger methods are needed to develop, review, and update the KB, as well as to regularly evaluate the accuracy of bot’s responses.
@@ 8. Conclusion
Building a system that addresses patients’ information needs by providing expert-verified information is an open problem in healthcare. To tackle this, we propose a novel solution—a LLM-powered experts-in-the-loop chatbot framework, that utilizes RAG over a custom KB to provide synchronous responses, and utilizes experts to provide verified responses asynchronously. Our in-the-wild study involving 49 participants (patients and attendants) and 6 experts (4 doctors and 2 coordinators) demonstrated the positive impact of this technological intervention. Patients not only trusted the chatbot’s response, but were also willing to wait, understanding that their busy doctors were verifying them. Simultaneously, doctors and coordinators, despite their hectic schedules, made time during breaks and at home to help patients by verifying LLM-generated responses. The favorable feedback from various stakeholders in the healthcare ecosystem indicates that a chatbot, delivered through ubiquitous smartphones and WhatsApp, can effectively enhance information access in critical healthcare settings, even for individuals with limited literacy and technology experience.
@@ Acknowledgements.
- Akerkar and Bichile (2004) Shashank M Akerkar and Lata S Bichile. 2004. Doctor patient relationship: changing dynamics in the information age. Journal of postgraduate medicine 50, 2 (2004), 120–122.
- Arora et al. (2008) Neeraj K Arora, Bradford W Hesse, Barbara K Rimer, Kasisomayajula Viswanath, Marla L Clayman, and Robert T Croyle. 2008. Frustrated and confused: the American public rates its cancer-related information-seeking experiences. Journal of general internal medicine 23 (2008), 223–228.
@@
- Beasley (2009) Robert E Beasley. 2009. Short message service (SMS) texting symbols: A functional analysis of 10,000 cellular phone text messages. The Reading Matrix 9, 2 (2009).
@@
- Bernstein et al. (2023) Isaac A Bernstein, Youchen Victor Zhang, Devendra Govil, Iyad Majid, Robert T Chang, Yang Sun, Ann Shue, Jonathan C Chou, Emily Schehlein, Karen L Christopher, et al. 2023. Comparison of ophthalmologist and large language model chatbot responses to online patient eye care questions. JAMA Network Open 6, 8 (2023), e2330320–e2330320.
- Bérubé et al. (2021) Caterina Bérubé, Zsolt Ferenc Kovacs, Elgar Fleisch, and Tobias Kowatsch. 2021. Reliability of Commercial Voice Assistants’ Responses to Health-Related Questions in Noncommunicable Disease Management: Factorial Experiment Assessing Response Rate and Source of Information. Journal of Medical Internet Research 23, 12 (2021), e32161.
- Bhat et al. (2021) Karthik S Bhat, Mohit Jain, and Neha Kumar. 2021. Infrastructuring Telehealth in (In)Formal Patient-Doctor Contexts. Proc. ACM Hum.-Comput. Interact. 5, CSCW2, Article 323 (oct 2021), 28 pages. https://doi.org/10.1145/3476064
@@
- Bodenheimer and Smith (2013) Thomas S Bodenheimer and Mark D Smith. 2013. Primary care: proposed solutions to the physician shortage without training more physicians. Health Affairs 32, 11 (2013), 1881–1886.
@@
- Chandwani and Kulkarni (2016) Rajesh Chandwani and Vaibhavi Kulkarni. 2016. Who’s the Doctor? Physicians’ Perception of Internet Informed Patients in India. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’16). Association for Computing Machinery, New York, NY, USA, 3091–3102. https://doi.org/10.1145/2858036.2858500
@@
- Chung (2017) Chia-Fang Chung. 2017. Supporting patient-provider communication and engagement with personal informatics data. In Proceedings of the 2017 ACM International Joint Conference on Pervasive and Ubiquitous Computing and Proceedings of the 2017 ACM International Symposium on Wearable Computers (Maui, Hawaii) (UbiComp ’17). Association for Computing Machinery, New York, NY, USA, 335–338. https://doi.org/10.1145/3123024.3123197
@@
- De Choudhury et al. (2014) Munmun De Choudhury, Meredith Ringel Morris, and Ryen W. White. 2014. Seeking and sharing health information online: comparing search engines and social media. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Toronto, Ontario, Canada) (CHI ’14). Association for Computing Machinery, New York, NY, USA, 1365–1376. https://doi.org/10.1145/2556288.2557214
- Denecke et al. (2024) Kerstin Denecke, Richard May, and Octavio Rivera Romero. 2024. Potential of Large Language Models in Health Care: Delphi Study. J Med Internet Res 26 (13 May 2024), e52399. https://doi.org/10.2196/52399
- Desai and Chin (2023) Smit Desai and Jessie Chin. 2023. OK Google, Let’s Learn: Using Voice User Interfaces for Informal Self-Regulated Learning of Health Topics among Younger and Older Adults. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Hamburg¡/city¿, ¡country¿Germany¡/country¿, ¡/conf-loc¿) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 847, 21 pages. https://doi.org/10.1145/3544548.3581507
- Ding et al. (2020) Xianghua Ding, Xinning Gui, Xiaojuan Ma, Zhaofei Ding, and Yunan Chen. 2020. Getting the Healthcare We Want: The Use of Online ”Ask the Doctor” Platforms in Practice. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3313831.3376699
@@
- Fochsen et al. (2006) Grethe Fochsen, Kirti Deshpande, and Anna Thorson. 2006. Power Imbalance and Consumerism in the Doctor-Patient Relationship: Health Care Providers’ Experiences of Patient Encounters in a Rural District in India. Qualitative Health Research 16, 9 (2006), 1236–1251. https://doi.org/10.1177/1049732306293776 arXiv:https://doi.org/10.1177/1049732306293776 PMID: 17038755.
- Fogel and Kvedar (2018) Alexander L Fogel and Joseph C Kvedar. 2018. Artificial intelligence powers digital medicine. NPJ digital medicine 1, 1 (2018), 5.
- Gordon et al. (2020) Howard S. Gordon, Pooja Solanki, Barbara G. Bokhour, and Ravi K. Gopal. 2020. “I’m Not Feeling Like I’m Part of the Conversation” Patients’ Perspectives on Communicating in Clinical Video Telehealth Visits. Journal of General Internal Medicine 35, 6 (01 Jun 2020), 1751–1758. https://doi.org/10.1007/s11606-020-05673-w
- Grönvall and Verdezoto (2013) Erik Grönvall and Nervo Verdezoto. 2013. Beyond self-monitoring: understanding non-functional aspects of home-based healthcare technology. In Proceedings of the 2013 ACM International Joint Conference on Pervasive and Ubiquitous Computing (Zurich, Switzerland) (UbiComp ’13). Association for Computing Machinery, New York, NY, USA, 587–596. https://doi.org/10.1145/2493432.2493495
- Henry et al. (2022) Katharine E Henry, Rachel Kornfield, Anirudh Sridharan, Robert C Linton, Catherine Groh, Tony Wang, Albert Wu, Bilge Mutlu, and Suchi Saria. 2022. Human–machine teaming is key to AI adoption: clinicians’ experiences with a deployed machine learning system. NPJ digital medicine 5, 1 (2022), 97.
- Holmes et al. (2023) Samuel Holmes, Raymond Bond, Anne Moorhead, Jane Zheng, Vivien Coates, and Michael McTear. 2023. Towards Validating a Chatbot Usability Scale. In Design, User Experience, and Usability, Aaron Marcus, Elizabeth Rosenzweig, and Marcelo M. Soares (Eds.). Springer Nature Switzerland, Cham, 321–339.
- Iliashenko et al. (2019) Oksana Iliashenko, Zilia Bikkulova, and Alissa Dubgorn. 2019. Opportunities and challenges of artificial intelligence in healthcare. In E3S Web of Conferences, Vol. 110. EDP Sciences, 02028.
- Ittarat et al. (2023) Mantapond Ittarat, Wisit Cheungpasitporn, and Sunee Chansangpetch. 2023. Personalized Care in Eye Health: Exploring Opportunities, Challenges, and the Road Ahead for Chatbots. Journal of Personalized Medicine 13, 12 (2023), 1679.
@@
- Kapania et al. (2022) Shivani Kapania, Oliver Siy, Gabe Clapper, Azhagu Meena SP, and Nithya Sambasivan. 2022. ”Because AI is 100% right and safe”: User Attitudes and Sources of AI Authority in India. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿New Orleans¡/city¿, ¡state¿LA¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 158, 18 pages. https://doi.org/10.1145/3491102.3517533
@@
- Kocaballi et al. (2020) A Baki Kocaballi, Kiran Ijaz, Liliana Laranjo, Juan C Quiroz, Dana Rezazadegan, Huong Ly Tong, Simon Willcock, Shlomo Berkovsky, and Enrico Coiera. 2020. Envisioning an artificial intelligence documentation assistant for future primary care consultations: A co-design study with general practitioners. Journal of the American Medical Informatics Association 27, 11 (08 2020), 1695–1704. https://doi.org/10.1093/jamia/ocaa131 arXiv:https://academic.oup.com/jamia/article-pdf/27/11/1695/34363907/ocaa131.pdf
@@
- Makarem and Antoun (2016) Nisrine N. Makarem and Jumana Antoun. 2016. Email communication in a developing country: different family physician and patient perspectives. Libyan Journal of Medicine 11, 1 (2016), 32679. https://doi.org/10.3402/ljm.v11.32679 arXiv:https://doi.org/10.3402/ljm.v11.32679 PMID: 28349842.
- Malki et al. (2024) Lisa Mekioussa Malki, Dilisha Patel, and Aneesha Singh. 2024. ”The Headline Was So Wild That I Had To Check”: An Exploration of Women’s Encounters With Health Misinformation on Social Media. Proc. ACM Hum.-Comput. Interact. 8, CSCW1, Article 128 (apr 2024), 26 pages. https://doi.org/10.1145/3637405
- Mamykina et al. (2008) Lena Mamykina, Elizabeth Mynatt, Patricia Davidson, and Daniel Greenblatt. 2008. MAHI: investigation of social scaffolding for reflective thinking in diabetes management. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Florence, Italy) (CHI ’08). Association for Computing Machinery, New York, NY, USA, 477–486. https://doi.org/10.1145/1357054.1357131
@@
- Office of the Registrar General & Census Commissioner (2011) Government of India Office of the Registrar General & Census Commissioner, India; Ministry of Home Affairs. 2011. Census of India 2011: C-16: Population by mother tongue, Karnataka - 2011. https://censusindia.gov.in/nada/index.php/catalog/10208/study-description Accessed: 2024-07-24.
@@
- OpenAI (2024) OpenAI. 2024. Usage Policies. https://openai.com/policies/usage-policies/ Accessed: 2024-06-18.
@@
- Osman Andersen et al. (2021) Tariq Osman Andersen, Francisco Nunes, Lauren Wilcox, Elizabeth Kaziunas, Stina Matthiesen, and Farah Magrabi. 2021. Realizing AI in Healthcare: Challenges Appearing in the Wild. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems (, Yokohama, Japan,) (CHI EA ’21). Association for Computing Machinery, New York, NY, USA, Article 108, 5 pages. https://doi.org/10.1145/3411763.3441347
- Perrier et al. (2015) Trevor Perrier, Nicola Dell, Brian DeRenzi, Richard Anderson, John Kinuthia, Jennifer Unger, and Grace John-Stewart. 2015. Engaging Pregnant Women in Kenya with a Hybrid Computer-Human SMS Communication System. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (Seoul, Republic of Korea) (CHI ’15). Association for Computing Machinery, New York, NY, USA, 1429–1438. https://doi.org/10.1145/2702123.2702124
- Pfeifer Vardoulakis et al. (2012) Laura Pfeifer Vardoulakis, Amy Karlson, Dan Morris, Greg Smith, Justin Gatewood, and Desney Tan. 2012. Using mobile phones to present medical information to hospital patients. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Austin, Texas, USA) (CHI ’12). Association for Computing Machinery, New York, NY, USA, 1411–1420. https://doi.org/10.1145/2207676.2208601
- Prasad et al. (2020) Manya Prasad, Sumit Malhotra, Mani Kalaivani, Praveen Vashist, and Sanjeev K Gupta. 2020. Gender differences in blindness, cataract blindness and cataract surgical coverage in India: a systematic review and meta-analysis. British Journal of Ophthalmology 104, 2 (2020), 220–224.
@@
- Rajashekar et al. (2024) Niroop Channa Rajashekar, Yeo Eun Shin, Yuan Pu, Sunny Chung, Kisung You, Mauro Giuffre, Colleen E Chan, Theo Saarinen, Allen Hsiao, Jasjeet Sekhon, Ambrose H Wong, Leigh V Evans, Rene F. Kizilcec, Loren Laine, Terika Mccall, and Dennis Shung. 2024. Human-Algorithmic Interaction Using a Large Language Model-Augmented Artificial Intelligence Clinical Decision Support System. In Proceedings of the CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Honolulu¡/city¿, ¡state¿HI¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 442, 20 pages. https://doi.org/10.1145/3613904.3642024
@@
- Seitz et al. (2022) Lennart Seitz, Sigrid Bekmeier-Feuerhahn, and Krutika Gohil. 2022. Can we trust a chatbot like a physician? A qualitative study on understanding the emergence of trust toward diagnostic chatbots. International Journal of Human-Computer Studies 165 (2022), 102848. https://doi.org/10.1016/j.ijhcs.2022.102848
@@
- Tan et al. (2024) Ting Fang Tan, Kabilan Elangovan, Liyuan Jin, Yao Jie, Li Yong, Joshua Lim, Stanley Poh, Wei Yan Ng, Daniel Lim, Yuhe Ke, et al. 2024. Fine-tuning Large Language Model (LLM) Artificial Intelligence Chatbots in Ophthalmology and LLM-based evaluation using GPT-4. arXiv preprint arXiv:2402.10083 (2024).
@@
- Thieme et al. (2023a) Anja Thieme, Maryann Hanratty, Maria Lyons, Jorge Palacios, Rita Faia Marques, Cecily Morrison, and Gavin Doherty. 2023a. Designing Human-centered AI for Mental Health: Developing Clinically Relevant Applications for Online CBT Treatment. ACM Trans. Comput.-Hum. Interact. 30, 2, Article 27 (mar 2023), 50 pages. https://doi.org/10.1145/3564752
- Thieme et al. (2023b) Anja Thieme, Aditya Nori, Marzyeh Ghassemi, Rishi Bommasani, Tariq Osman Andersen, and Ewa Luger. 2023b. Foundation Models in Healthcare: Opportunities, Risks & Strategies Forward. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems (, Hamburg, Germany,) (CHI EA ’23). Association for Computing Machinery, New York, NY, USA, Article 512, 4 pages. https://doi.org/10.1145/3544549.3583177
@@
- Tuli et al. (2018) Anupriya Tuli, Shaan Chopra, Neha Kumar, and Pushpendra Singh. 2018. Learning from and with Menstrupedia: Towards Menstrual Health Education in India. Proc. ACM Hum.-Comput. Interact. 2, CSCW, Article 174 (nov 2018), 20 pages. https://doi.org/10.1145/3274443
@@
- Wei et al. (2024) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H. Chi, Quoc V. Le, and Denny Zhou. 2024. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22). Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages.
@@
- Wutz et al. (2023) Maximilian Wutz, Marius Hermes, Vera Winter, and Juliane Köberlein-Neu. 2023. Factors influencing the acceptability, acceptance, and adoption of conversational agents in health care: integrative review. Journal of Medical Internet Research 25 (2023), e46548.
@@
- Yang et al. (2024) Ziqi Yang, Xuhai Xu, Bingsheng Yao, Ethan Rogers, Shao Zhang, Stephen Intille, Nawar Shara, Guodong Gordon Gao, and Dakuo Wang. 2024. Talk2Care: An LLM-based Voice Assistant for Communication between Healthcare Providers and Older Adults. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 2, Article 73 (may 2024), 35 pages. https://doi.org/10.1145/3659625
- Ye et al. (2020) Qunru Ye, Yanxian Chen, William Yan, Wei Wang, Jingxian Zhong, Cong Tang, Andreas Müller, Bo Qiu, et al. 2020. Female gender remains a significant barrier to access cataract surgery in South Asia: A systematic review and meta-analysis. Journal of ophthalmology 2020 (2020).
- Yeh et al. (2022) Su-Fang Yeh, Meng-Hsin Wu, Tze-Yu Chen, Yen-Chun Lin, XiJing Chang, You-Hsuan Chiang, and Yung-Ju Chang. 2022. How to Guide Task-oriented Chatbot Users, and When: A Mixed-methods Study of Combinations of Chatbot Guidance Types and Timings. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 488, 16 pages. https://doi.org/10.1145/3491102.3501941
- Yilmaz and Dogan (2024) Ibrahim Edhem Yilmaz and Levent Dogan. 2024. Talking technology: exploring chatbots as a tool for cataract patient education. Clinical and Experimental Optometry (2024), 1–9.
@@
- Zhang (2010) Yan Zhang. 2010. Contextualizing consumer health information searching: an analysis of questions in a social Q&A community. In Proceedings of the 1st ACM international health informatics symposium. 210–219.